# Beyond Words: On Large Language Models Actionability in Mission-Critical Risk Analysis

### v4

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 18th ACM/IEEE International Symposium on Empirical Software Engineering and Measurement; October 20–25, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Risk managementccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software design tradeoffs
Laydown Abstract
@@ 2.1. Definitions
2.2. Retrieve-Augmented Generation
@@ 4. Methodology
Based on our goal, we defined the following three Research Questions ($RQ_{s}$).
RQ1 Do LLMs generate accurate RA?
@@
However, RA is not just about generating reports; it must be actionable. Therefore, we ask:
RQ2 Do LLMs generate actionable and comprehensive RA?
Actionable and comprehensive insights from LLM conducting risk analysis are essential in decision-making processes across various domains. Whether it is assessing financial risks, cybersecurity threats, or environmental hazards, the insights an LLM provides must translate into tangible actions (Ge et al., 2023). By transforming complex scenario analysis into clear recommendations, actionable output allows practitioners to implement preventive measures, allocate resources efficiently, and mitigate potential risks proactively. Risk analysis remains purely academic without actionable insights, failing to safeguard against emerging threats (Aven, 2015).
@@
Finally, LLM has the advantage of analyzing vast amounts of data in less time than humans and storing new information on demand via RAG. Therefore, we ask:
RQ3 Do LLMs outperform human experts in RA?
The level of human expertise and years spent in the field directly impact PSRA quality and proficiency. We can measure the time required for a new team member to attain the proficiency level of a senior team member in years or even decades. In contrast, LLMs can rapidly assimilate years of training within mere minutes. Therefore, it is essential to investigate the performance of the LLM model in terms of accuracy and comprehensiveness with respect to the HEs. Table 1 provides an overview of the experts’ profiles for reviewers and test participants.
@@ 4.3. Data Analysis
To answer $RQ_{2}$, we measured the actionability and comprehensiveness of the model’s RA. Although GT comes from finalized RAs, we considered the possibility for human experts to overlook some details, thus failing to identify specific or more subtle threats. We call this threat “hidden risks”. Therefore, we also evaluated the ability of models and human experts to extend the ground truth (GT) to measure if the finalized RAs (GT) were spotless or if hidden risks were present. Table 3 presents the metrics we used to evaluate the scenarios. More specifically, we instructed the HRs to analyze the HE and the model’s output according to Table LABEL:tab:prediction_types and Table 3.
@@ 6.2. Model Explainability
In our context, according to Arrieta et al. (2020) definition, an actionable output is self-explanatory regarding the model’s inner workings. Specifically, in the RAG-assisted model, the knowledge of acronyms, threats, and vulnerabilities is traceable back to the laws. For instance, the traceability is guaranteed by the assistant API [^4] https://platform.openai.com/docs/assistants/overview responds containing annotations that explain what information was taken from which document.
@@ 6.4. Comparison with Human Experts
According to Table 10, we noted that HEs never extended the GT. Therefore, we can think that specific scenarios hide risks humans cannot grasp.

### v5

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement; October 24–25, 2024; Barcelona, Spainbooktitle: Proceedings of the 18th ACM / IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM ’24), October 24–25, 2024, Barcelona, Spaindoi: 10.1145/3674805.3695401isbn: 979-8-4007-1047-6/24/10ccs: Software and its engineering Risk managementccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software design tradeoffs
Lay Abstract
@@ 2.1. Definitions
2.2. Retrieval-Augmented Generation
@@ 4. Methodology
[…]
@@
[…]
@@
[…]
@@ 4.3. Data Analysis
To answer $RQ_{2}$, we measured the actionability and comprehensiveness of the model’s RA. Although GT comes from finalized RAs, we considered the possibility for human experts to overlook some details, thus failing to identify specific or more subtle threats. We call this threat “hidden risks”. Therefore, we also evaluated the ability of models and human experts to extend the ground truth (GT) to measure if the finalized RAs (GT) were spotless or if hidden risks were present. Table 3 presents the metrics we used to evaluate the scenarios. More specifically, we instructed the HRs to analyze the HE and the model’s output according to Table 3.
@@ 6.1. Model Accuracy
1. LLMs as RA copilot.
LLMs proved to be quicker, more comprehensive, and actionable than humans with a fraction of the time spent. Therefore, considering a trade-off between accuracy and speed, LLMs are a valuable tool in the hands of experts to quickly asses complex scenarios.
[…]
@@ 6.2. Model Explainability
[…]
In a mission-critical risk analysis context, we can use actionability as a proxy metric for measuring model explainability.
@@ 6.3. LLMs Hallucination
3. Hallucination and human performance.
RAG-assisted LLMs reduce hallucinations and can challenge human experts in describing vulnerabilities.
[…]
@@ 6.4. Comparison with Human Experts
[…]
In a mission-critical risk analysis context, we should assess each possible risk to the best of our knowledge. Nonetheless, humans can overlook details that LLMs can track down.
@@
[…]
5. To RAG or Fine-Tune, a bards dilemma.
Choosing the suitable model is essential to ensure hitting the right target. According to our findings, FTM proved to be the most accurate model, while RAG is unmatched in discovering hidden risks and hallucinating the less. In contrast, the base model excels in actionability and comprehensiveness.
@@ 10. Conclusions
Acknowledgements.
This work has been partially funded by Business Finland project 6GBridge/6GSoft.